### v1

# Entanglement growth and information capacity in a quasiperiodic system with a single-particle mobility edge

Abstract
We investigate the quantum dynamics of a one-dimensional quasiperiodic system featuring a single-particle mobility edge (SPME), described by the generalized Aubry-André (GAA) model. This model offers a unique platform to study the consequences of coexisting localized and extended eigenstates, which contrasts sharply with the abrupt localization transition in the standard Aubry-André model. We analyze the system’s response to a quantum quench through two complementary probes: entanglement entropy (EE) and subsystem information capacity (SIC). We find that the SPME induces a smooth crossover in all dynamical signatures. The EE saturation value exhibits a persistent volume-law scaling in the mobility-edge phase, with an entropy density that continuously decreases as the number of available extended states decreases. Complementing this, the SIC profile interpolates between the linear ramp characteristic of extended systems and the information trapping behavior of localized ones, directly visualizing the mixed nature of the underlying spectrum. Our results establish unambiguous dynamical fingerprints of a mobility edge, providing a crucial non-interacting benchmark for understanding information and entanglement dynamics in more complex systems with mixed phases.
Introduction.— Understanding thermalization and its breakdown in isolated quantum systems is a fundamental problem in quantum many-body physics D’Alessio et al. (2016); Deutsch (2018). While generic interacting systems typically thermalize following the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994); Rigol et al. (2008), this behavior can be subverted by strong quenched disorder. For example, quenched disorder can cause Anderson localization in non-interacting systems Anderson (1958) and many-body localization (MBL) can emerge when interaction is turned on, leading to a nonergodic phase where memory of the initial state persists Fleishman and Anderson (1980); Basko et al. (2006); Nandkishore and Huse (2015); Abanin et al. (2019).
Besides random disorder, quasiperiodic potentials offer an alternative path to localization Iyer et al. (2013); Lee et al. (2017); Chandran and Laumann (2017); Setiawan et al. (2017); Schreiber et al. (2015); Khemani et al. (2017); Zhang and Yao (2018, 2019); Wang et al. (2020); Zhai et al. (2020). The standard Aubry-André (AA) model, for instance, shows a sharp transition where all single-particle states become localized at the critical potential strength Aubry and André (1980); Harper (1955). A more intriguing scenario is presented by the generalized Aubry-André (GAA) model Thouless (1983); Han et al. (1994); Biddle and Das Sarma (2010); Ganeshan et al. (2015), which holds a more complex spectral structure with a single-particle mobility edge (SPME)—an energy threshold separating coexisting localized and extended eigenstates. The recent experimental realization of such models in ultracold atoms Kohlert et al. (2019) and photonic lattices Verbin et al. (2015); Lüschen et al. (2018) have provided excellent platforms for theoretical study. Understanding the precise dynamical signatures in the non-interacting setting with SPME is a critical first step, providing an essential baseline for tackling the more complex and debated questions of thermalization in interacting systems that may host a mobility edge Li et al. (2015); Modak and Mukerjee (2015); Deng et al. (2017); Li et al. (2016); Modak et al. (2018); Huang et al. (2023).
Concepts from quantum information theory offer powerful tools for probing quantum many-body systems Amico et al. (2008); Eisert et al. (2010); Abanin et al. (2019). Among these, entanglement entropy (EE) is a primary probe of correlations and quantum phase structures. The scaling of EE with subsystem size clearly separates different quantum phases, for example, ground states of gapped systems follow an area law. In terms of time-evolved states, Anderson localization systems follow an area law, while thermalizing systems follow a volume law Bauer and Nayak (2013); Kaufman et al. (2016). Beyond its late-time saturation value, the initial growth of entanglement entropy (EE) after a quantum quench serves as a powerful dynamical probe. Linear growth is a hallmark of chaotic thermalizing systems, logarithmic growth is characteristic of many-body localization (MBL), and a near-complete lack of growth signifies Anderson localization Bardarson et al. (2012); Nanduri et al. (2014); Lewis-Swan et al. (2019). This raises a key question for the GAA model: How does entanglement evolve in a system defined by the coexistence of localized and extended modes? Does it obey a volume law due to the extended states, an area law governed by the localized ones, or does a new, intermediate behavior emerge?
While EE quantifies the overall magnitude of entanglement, it does not fully resolve how quantum information propagates and where it is stored. Other tools reveal different aspects of information dynamics. For example, mutual information tracks shared correlations Wolf et al. (2008), and its extensions like tripartite mutual information can detect delocalized entanglement and information scrambling Hosur et al. (2016). Moreover, out-of-time-ordered correlators are frequently used to diagnose the spreading of local information into non-local correlations—a signature of quantum chaos Maldacena et al. (2016); Fan et al. (2017); Nahum et al. (2018). However, to obtain a more direct, spatially resolved picture of how information propagates and is retained in space-time, we employ the subsystem information capacity the subsystem information capacity (SIC) Chen et al. (2024), a recently introduced probe of information propagation and retention. This quantity investigates the information flow by viewing a subsystem’s evolution as an effective quantum channel whose capacity is shaped by the system’s overall dynamics Holevo (2012). Defined as the mutual information $I(A{:}R)$ between an output subsystem $A$ and a reference system $R$ initially entangled with an input region $E$, SIC measures how much of the initial quantum information can be recovered from $A$ after some time. This measure is closely related to the quantum coherent information of the effective channel $\mathcal{E}_{E\to A}$ Schumacher and Nielsen (1996); Chen et al. (2024), reflecting the system’s ability to preserve quantum information and acting as a measure of the channel’s single-shot quantum capacity Lloyd (1997); Devetak and Winter (2005). Importantly, SIC shows distinct spatial profiles in different dynamical phases—from information trapping behavior in localized systems to linear growth in extended ones Chen et al. (2024). This makes SIC an ideal diagnostic for the mixed dynamics expected in the GAA model, allowing us to ask whether the contributions of localized and extended modes can be spatially separated.
In this work, we employ two complementary quantum information probes—entanglement entropy and subsystem information capacity—to conduct a comprehensive analysis of the quench dynamics in the non-interacting generalized Aubry-André (GAA) model. By contrasting its behavior with the abrupt localization transition in the standard AA model, we establish clear and unambiguous dynamical fingerprints of an SPME. We find that the SPME induces a crossover in both dynamical signatures, replacing the sharp transition. Specifically, the EE saturation value exhibits a persistent volume-law scaling, yet with a continuously tunable entropy density controlled by the number of available extended states. Complementing this, the SIC profile directly visualizes the mixed spectral nature, interpolating between the linear ramp characteristic of extended systems and the information trapping of localized ones. Our results provide a definitive non-interacting benchmark crucial for understanding information and entanglement dynamics in interacting systems with mixed phases.
Model and Methods.— We study the dynamics of noninteracting spinless fermions on a one-dimensional lattice with open boundary conditions, described by the Generalized Aubry-André (GAA) Hamiltonian
$H=-t\sum_{i=1}^{L-1}(c_{i}^{\dagger}c_{i+1}+\text{h.c.})+\sum_{i=1}^{L}\mu_{i}\,c_{i}^{\dagger}c_{i},$ (1)
where $c_{i}^{\dagger}$ ($c_{i}$) creates (annihilates) a fermion at site $i$, $L$ is the system size, and $t$ is the nearest-neighbor hopping, set to unity ($t=1$). The on-site quasiperiodic potential is
$\mu_{i}=2\lambda\frac{\cos(2\pi bi+\phi)}{1-a\cos(2\pi bi+\phi)},$ (2)
[…]
The standard AA model ($a=0$) features a sharp localization transition at $\lambda=t$, where all single-particle eigenstates abruptly transition from being extended to localized Aubry and André (1980). For $a\neq 0$, the GAA model hosts a single-particle mobility edge at energy $E_{c}$, given by Ganeshan et al. (2015):
$aE_{c}=2\,\mathrm{sgn}(\lambda)(|t|-|\lambda|).$ (3)
This critical energy leads to the phase diagram in Fig. 1, showing a completely extended, a completely localized, and an intermediate phase with a mobility edge where extended and localized states coexist.
Figure 1: Phase diagram of the GAA model, featuring three distinct dynamical regimes based on the potential strength $\lambda/|t|$ and deformation $a$. The model has completely extended (pink) and localized (green) phases, separated by an intermediate phase (yellow) where extended and localized states coexist due to the SPME.
The key feature of the GAA model for $a\neq 0$ is the coexistence of localized and extended states. The effect of the SPME is visualized by the inverse participation ratio (IPR) of the eigenstates,
$\mathrm{IPR}_{n}=\frac{\sum_{i=1}^{L}|\psi_{n,i}|^{4}}{(\sum_{i=1}^{L}|\psi_{n,i}|^{2})^{2}},$ (4)
where $\psi_{n}$ is the $n$-th eigenstate of the system. As shown in Fig. 2 for $a=0.3$, the eigenstates are clearly separated into extended (low IPR) and localized (high IPR) modes by the energy-dependent SPME. In the following, we study the quench dynamics governed by this Hamiltonian to explore the consequences of this mixed spectral structure.
Figure 2: Energy spectrum of the GAA model for $a=0.3$ and $L=200$, with eigenstates colored by their IPR. The mobility edge $E_{c}$ (red lines), from Eq. (3), clearly separates extended states (low IPR, bright cyan) from localized states (high IPR, dark teal/black). This energy-dependent separation is a key feature of the SPME phase.
Entanglement Dynamics.— We consider unitary evolution $|\Psi(t)\rangle=e^{-iHt}|\Psi(0)\rangle$ after a quantum quench from an unentangled product state at half filling ($N=L/2$). We use the Néel state ($|1010\dots\rangle$) as the initial state throughout the main text (see SM for results from different initial states). To measure the dynamics, we compute the half-chain entanglement entropy (EE), $S(t)=-\mathrm{Tr}[\rho_{L/2}(t)\ln\rho_{L/2}(t)]$, where $\rho_{L/2}(t)$ is the reduced density matrix of the system’s left half.
First, we examine the early-time EE growth velocity, $v_{S}$, extracted from the initial linear growth regime (see SM for entanglement dynamics plots). As shown in Fig. 3, for the standard AA model ($a=0$), $v_{S}$ decreases monotonically with increasing $\lambda$. For $a>0$, the behavior is more complex: while $v_{S}$ still decreases with $\lambda$, distinct plateaus appear. These plateaus are a non-trivial signature of the energy spectrum of the GAA model, appearing when the SPME aligns with the system’s energy gaps (see Fig.2). Notably, the growth speed of EE $v_{S}$ begins to decrease as long as $\lambda>0$ even in the fully extended phases. This reflects the fact that any quasiperiodic potential, no matter how weak, slows down particle propagation. This early-time character is in stark contrast to the behavior of late-time EE $S_{\mathrm{sat}}$ introduced next, which remains constant at fully extended phase and begins to drop at the intermediate phase.
Figure 3: Early-time EE growth velocity $v_{S}$ versus $\lambda/|t|$ for different $a$. The initial state is a Néel state with $L=200$. While the AA model ($a=0$) shows a simple decrease, the GAA model ($a>0$) exhibits plateaus. These features are signatures of the energy gaps in the SPME phase.
Next, we analyze the long-time saturation EE, $S_{\mathrm{sat}}$, extracted by averaging the EE over a long-time window after it reaches a steady value. As shown in Fig. 4, for the standard AA model ($a=0$), $S_{\mathrm{sat}}$ drops sharply at the $\lambda=t$ transition, signaling the phase transition nature from a volume-law to an area-law phase. In contrast, the GAA models ($a>0$) exhibit a smooth crossover, indicating a gradual suppression of entanglement generation rather than a complete halt. The visible plateaus again correspond to the SPME traversing the spectral gaps.
Figure 4: Saturation EE $S_{\mathrm{sat}}$ versus $\lambda/|t|$ after a quench from the Néel state for $L=200$. The AA model ($a=0$, blue) has a sharp drop at the $\lambda=t$ transition. The GAA models ($a>0$) show a smooth crossover, indicating partial delocalization caused by the SPME.
[…]
Figure 5: Finite-size scaling of the saturation entropy density $S_{\mathrm{sat}}/(L/2)$ versus inverse system size $1/L$. Each panel is for a different $a$. The data distinguishes area-law (density vanishes as $1/L\rightarrow 0$) from volume-law (density converges to a finite value) scaling. The persistent volume law for $a>0$ at intermediate $\lambda$ confirms the delocalizing effect of the SPME.
The underlying physical mechanism is revealed in Fig. 6, which plots $S_{\mathrm{sat}}$ against the number of extended states, $N_{e}$ (states with $|E|<|E_{c}|$). The two quantities are strongly correlated. This striking one-to-one correspondence demonstrates that the system’s capacity to generate volume-law entanglement is directly controlled by the fraction of available extended modes, which act as resources for thermalization. In other words, the decrease in $S_{\mathrm{sat}}$ tracks the reduction in available extended modes as $\lambda$ increases.
Figure 6: Connection between saturation entanglement $S_{\mathrm{sat}}$ (blue circles, left axis) and the number of extended states $N_{e}$ (green triangles, right axis) versus $\lambda/|t|$ for $a=0.3$ and $L=200$. $N_{e}$ is the number of states with $|E|<|E_{c}|$. The strong correlation shows that the system’s capability to develop volume-law entanglement is directly controlled by the number of available extended modes.
Figure 7: Steady-state SIC profile $I(A:R)$ versus subsystem size $|A|$ for $L=200$. The reference qubit is entangled with the center site $E=L/2$. (Top) The AA model ($a=0$) shows a sharp transition from a linear ramp to a step function. (Middle, Bottom) The GAA model ($a>0$) shows a smooth crossover. The hybrid nature of the profiles in the SPME phase—an initial jump followed by a slow ramp—is a direct visualization of the coexisting localized (information trapping) and extended (information spreading) modes.
Subsystem Information Capacity.— To move beyond the single scalar value of EE and gain a spatially resolved picture of information dynamics, we analyze the SIC Chen et al. (2024). This probe measures how quantum information, initially localized at a single site, propagates through the system. We prepare the system with a reference qubit $R$ maximally entangled with a single site $E$ of the chain:
$|\Phi\rangle=\frac{1}{\sqrt{2}}(|1\rangle_{E}|0\rangle_{R}+|0\rangle_{E}|1\rangle_{R}),$ (5)
with the rest of the chain in a product state. We compute the SIC between $R$ and a subsystem $A$ of size $|A|$ centered on $E=L/2$ after the time evolution, defined as
$I(A:R)=S(A)+S(R)-S(AR),$ (6)
where $S(\cdot)$ is the von Neumann entropy. Note that we use logarithmic base $2$ for the entropy convention in the SIC definition so that $I=2$ indicates the full information recovery capability. The spatial profile of steady-state $I(A:R)$ measures how information spreads: a linear ramp signifies ballistic transport in extended systems, whereas a step-function profile indicates information trapping in localized systems Chen et al. (2024).
The steady-state SIC profile (Fig. 7) provides a clear picture of the information distribution at late times. For the standard AA model ($a=0$), the profile shows a sharp transition from a perfect linear ramp ($\lambda/|t|<1$) to a sharp step function ($\lambda/|t|>1$). In contrast, the GAA model ($a>0$) shows a more smooth crossover. In the intermediate SPME regime, the SIC profiles are a hybrid of both behaviors: they rise steeply for small $|A|$, a signature of the localized modes trapping a portion of the information, and then continue to increase slowly for larger $|A|$, driven by the ballistic spreading through the remaining extended modes. This interpolation between a linear ramp and information trapping is a direct and powerful dynamical signature of the SPME.
To quantify the mixed behavior, we define $\text{SIC}_{\text{jump}}$ as the SIC value for a small subsystem size (we use $|A|=5$) intended to capture the information confined by localized modes due to the trapping nature. Fig. 8 plots $\text{SIC}_{\text{jump}}$ and the number of localized states, $N_{l}$ (states with $|E|>|E_{c}|$). For small $\lambda$ in the fully extended phase, $\text{SIC}_{\text{jump}}$ is near zero. As $\lambda$ increases in the intermediate phase with SPME, $\text{SIC}_{\text{jump}}$ rises sharply, tracking the increase in $N_{l}$. This powerful correlation between $\text{SIC}_{\text{jump}}$ and the number of localized states $N_{l}$ confirms the physical origin of the mixed SIC profile: the initial jump is a direct consequence of the growing fraction of localized states, which trap a corresponding amount of information around the initial site. We also present the late-time SIC profiles when the reference site is coupled to the boundary site in the SM, which further corroborates our findings.
Figure 8: Connection between the initial SIC jump, $\text{SIC}_{\text{jump}}$ (blue circles, left axis), and the number of localized states $N_{l}$ (green triangles, right axis) for the GAA model with $a=0.3$ and $L=200$. $N_{l}$ is the number of states with $|E|>|E_{c}|$. The close connection shows that the emergence of information confinement, captured in $\text{SIC}_{\text{jump}}$, is directly caused by the increase in localized states.
Table 1: Comparison of dynamical signatures in the standard AA and GAA models within SPME phases.
Probe Standard AA GAA
($a=0$) ($a>0$)
EE ($S_{\text{sat}}$) Sharp transition: SPME phase:
Volume $\to$ Area Persistent volume-law
SIC Profile Sharp transition: SPME phase:
ramp $\to$ step-func. Hybrid jump+ramp
Discussions and conclusion.— Our comprehensive analysis shows a clear difference between the dynamics of the standard AA model and the GAA model with an SPME. In the AA model ($a=0$), the transition is sudden: $S_{\mathrm{sat}}$ switches from volume-law to area-law scaling abruptly, and the SIC profile changes from a ramp to a step. In contrast, the GAA model ($a\neq 0$) shows a smooth crossover. This is characterized by intermediate $S_{\mathrm{sat}}$ values, mixed SIC profiles, and a persistent volume law with a decreasing entropy density. This smoothing is a direct result of the SPME, which allows localized and extended modes to coexist and control the balance between information spreading and confinement. Our main findings are summarized in Table 1. Our results provide clear, non-interacting dynamical signatures of a mobility edge, establishing an important reference for understanding thermalization in more complex quasiperiodic systems. These insights are relevant to the current debate on MBL in the presence of an SPME Li et al. (2015); Modak and Mukerjee (2015); Deng et al. (2017); Huang et al. (2023). While we focus on the paradigmatic GAA model, our findings on the dynamical signatures of coexisting states are expected to provide insights for other systems hosting mobility edges, such as those with long-range interactions or in higher dimensions Li et al. (2016); Zhang and Yao (2018); Huang et al. (2023).
In conclusion, we have studied the quantum dynamics in the generalized Aubry-André model, a paradigmatic system with a single-particle mobility edge. By leveraging the complementary power of entanglement entropy and subsystem information capacity, we have demonstrated that the SPME fundamentally reshapes the system’s dynamical response, converting the sharp localization transition into a smooth crossover. Instead of the sudden transition in the standard AA model, the SPME facilitates a regime with smoothly varying entanglement, hybrid information-spreading profiles, and a robust volume law with tunable entropy density. Our findings highlight that these quantum information probes are highly sensitive to a system’s spectral structure and that mobility edges reshape dynamical phase boundaries. This work provides a clear and crucial non-interacting reference for localization dynamics in quasiperiodic systems and paves the way for future studies of interacting systems, where the interplay between SPMEs and many-body effects remains a major open question. Furthermore, understanding the control over information flow offered by an SPME, from ballistic transport to perfect trapping, holds potential relevance for quantum technologies.
Acknowledgements. SXZ acknowledges the support from a start-up grant at IOP-CAS.
Data availability. Code implementation and numerical data for this manuscript are publicly accessible in Ref.  [^1] https://gitee.com/yuqiqing/sic_in_-gaa_model.
Supplemental Material for "Entanglement growth and information capacity in a quasiperiodic system with a single-particle mobility edge"
June 22, 2025
I Time evolution for entanglement entropy
In this section, we provide details on the extraction of the early-time entanglement growth velocity $v_{S}$ and the late-time saturation value $S_{\mathrm{sat}}$ from the time evolution of the half-chain entanglement entropy (EE). Figure S1 shows the time evolution of EE for a quench from the Néel state. The dynamics can be separated into two distinct regimes:
  1. 1.
    Early-time linear growth regime: For short times after the quench, the EE grows approximately linearly. We extract the velocity $v_{S}$ by performing a linear fit to the EE curve in this initial period. In practice, the time window $t\in[0,20]$ serves as a reasonable choice to identify the early-time regime in the system under investigation.
  2. 2.
    Late-time saturation regime: At long times, the EE stops growing and oscillates around a steady-state value. We calculate the saturation value $S_{\mathrm{sat}}$ by averaging the EE over a long-time window in this saturation regime, after the initial growth has ceased. In our calculation, we first quench the system for a sufficiently long time $t=10000$, and then take 1000 samples with an average time interval of $\langle dt\rangle=10$. To ensure adequate sampling along the EE profile, the interval $dt$ is randomly varied around the average.
The values of $v_{S}$ and $S_{\mathrm{sat}}$ presented in the main text and this supplemental material are obtained using this procedure.
Figure S1: Time evolution of half-chain entanglement entropy (EE) from the Néel state. The plots show EE as a function of time $t$ for different values of the potential strength $\lambda/|t|$ (from $\lambda/|t|=0$, top curve, to $\lambda/|t|=2.0$, bottom curve) and deformation parameter $a$. The system size is $L=200$. The dynamics clearly show an initial growth phase followed by the saturation stage.
II Entanglement growth from other initial states
To ensure the generality of our conclusions, we supplement the results from the Néel state with data from two other types of initial product states at half-filling: the domain-wall state ($|11\dots 100\dots 0\rangle$) and random product states. For the random states, we average the results over different random configurations of $L/2$ fermions.
Figure S2 shows the time evolution of EE for these states. While the qualitative behavior is similar to the Néel state, we note that for the domain-wall state, the oscillation period in the saturation regime is significantly longer. The saturation values (Fig. S3) also show trends consistent with the main text: a sharp transition for the AA model ($a=0$) and a smooth crossover for the GAA model ($a>0$). Notably, Fig. S4 demonstrates that for both the domain-wall and random initial states, the saturation EE $S_{\mathrm{sat}}$ is strongly correlated with the number of extended states $N_{e}$, reinforcing the universality of our central conclusion.
Figure S2: Time evolution of EE for other initial states. (Top row) Quench from a domain-wall state. (Bottom row) Quench from random product states (averaged). The parameters are $L=200$. The qualitative features of EE growth and saturation are consistent across different initial states.
Figure S3: Saturation EE for other initial states. (a) Domain-wall initial state. (b) Random initial states (averaged). The plots show $S_{\mathrm{sat}}$ versus $\lambda/|t|$. The smooth crossover in the GAA model ($a>0$) with plateaus corresponding to the spectrum gap is a robust feature.
Figure S4: Connection between saturation EE and the number of extended states for other initial states. (a) Domain-wall initial state. (b) Random initial states (averaged). Both plots are for $a=0.3$ and $L=200$. $S_{\mathrm{sat}}$ (blue circles, left axis) strongly correlates with $N_{e}$ (green triangles, right axis), confirming the conclusion from the main text.
III Steady-state SIC profile for edge-coupling and other initial states
To study boundary effects, we also investigate the steady-state SIC profile by coupling the reference site to an edge site ($E=1$). As shown in Fig. S5, the SIC profile for this edge-coupling setup is different from the center-coupling case, especially for the standard AA model. In this case, the SIC profile fails to act as a clear probe for the phase transition, as the profile varies smoothly with increasing $\lambda$ even in the extended phase. This failure might be due to the asymmetric information propagation with boundary reflections and interference.
Figure S5: Steady-state SIC profile $I(A:R)$ for the edge-coupling case. The initial state is Néel state. The reference qubit is entangled with the edge site $E=1$, and the subsystem $A$ starts from this edge.
Finally, we present the subsystem information capacity (SIC) results for different initial states mentioned before. Figures S6 and S7 show the SIC profiles for center- and edge-coupling, respectively. The results confirm that the hybrid SIC profile is a generic dynamical signature of the SPME, independent of the specific initial product state.
Figure S6: SIC profiles for center-coupling with other initial states. The reference qubit is entangled with the center site $E=L/2$. (Top row) Domain-wall initial state. (Bottom row) Random initial states (averaged). System size $L=200$, subsystem size $|A|$ is measured for the central region. The mixed SIC profile in the GAA model is a generic feature.
Figure S7: SIC profiles for edge-coupling with other initial states. The reference qubit is entangled with the edge site $E=1$. (Top row) Domain-wall initial state. (Bottom row) Random initial states (averaged). System size $L=200$. The results are consistent with those in the main text.