### v2

# Entanglement growth and information capacity in a quasiperiodic system with a single-particle mobility edge

Abstract
We investigate the quantum dynamics of a one-dimensional quasiperiodic system featuring a single-particle mobility edge (SPME), described by the generalized Aubry-André model. This model offers a unique platform to study the consequences of coexisting localized and extended eigenstates, which contrasts sharply with the abrupt localization transition in the standard Aubry-André model. We analyze the system’s response to a quantum quench through two complementary probes, entanglement entropy (EE) and subsystem information capacity (SIC). We find that the SPME induces a smooth crossover in all dynamical signatures. The EE saturation value exhibits a persistent volume-law scaling in the mobility-edge phase, with an entropy that continuously decreases as the number of available extended states decreases. Complementing this, the SIC profile interpolates between the linear ramp characteristic of extended systems and the information trapping behavior of localized ones, directly visualizing the mixed nature of the underlying spectrum. Our results establish unambiguous dynamical fingerprints of a mobility edge, providing a crucial non-interacting benchmark for understanding information and entanglement dynamics in more complex systems with mixed phases.
Introduction.— Understanding thermalization and its breakdown in isolated quantum systems is a fundamental problem in quantum many-body physics D’Alessio et al. (2016); Deutsch (2018). While generic interacting systems typically thermalize following the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994); Rigol et al. (2008), this behavior can be subverted by strong quenched disorder. For example, quenched disorder can cause Anderson localization in non-interacting systems Anderson (1958) and many-body localization (MBL) can emerge when interaction is turned on, leading to a nonergodic phase where memory of the initial state persists Fleishman and Anderson (1980); Basko et al. (2006); Nandkishore and Huse (2015); Abanin et al. (2019).
Besides random disorder, quasiperiodic potentials Thouless (1983); Han et al. (1994); Iyer et al. (2013); Schreiber et al. (2015); Lee et al. (2017); Khemani et al. (2017); Chandran and Laumann (2017); Setiawan et al. (2017); Zhang and Yao (2018, 2019); Wang et al. (2020); Zhai et al. (2020) offer an alternative path to localization. The standard Aubry-André (AA) model, for instance, shows a sharp transition where all single-particle states become localized at the critical potential strength Aubry and André (1980); Harper (1955). A more intriguing scenario is presented by the generalized Aubry-André (GAA) model Biddle and Das Sarma (2010); Ganeshan et al. (2015), which holds a more complex spectral structure with a single-particle mobility edge (SPME)—an energy threshold separating coexisting localized and extended eigenstates. The recent experimental realization of such models in ultracold atoms Lüschen et al. (2018); Kohlert et al. (2019) and photonic lattices Verbin et al. (2015) have provided excellent platforms for theoretical study. Understanding the precise dynamical signatures in the non-interacting setting with SPME is a critical first step, providing an essential baseline for tackling the more complex and debated questions of thermalization in interacting systems that may host a mobility edge Li et al. (2015); Modak and Mukerjee (2015); Li et al. (2016); Deng et al. (2017); Modak et al. (2018); Huang et al. (2023).
Concepts from quantum information theory offer powerful tools for probing quantum many-body systems Amico et al. (2008); Eisert et al. (2010); Abanin et al. (2019). Among these, entanglement entropy (EE) is a primary probe of correlations and quantum phase structures. The scaling of EE with subsystem size clearly separates different quantum phases, for example, ground states of gapped systems follow an area law. In terms of time-evolved states, Anderson localization systems follow an area law, while thermalizing systems follow a volume law Bauer and Nayak (2013); Kaufman et al. (2016). Beyond its late-time saturation value, the initial growth of EE after a quantum quench serves as a powerful dynamical probe. Linear growth is a hallmark of chaotic thermalizing systems, logarithmic growth is characteristic of MBL, and a near-complete lack of growth signifies Anderson localization Bardarson et al. (2012); Nanduri et al. (2014); Lewis-Swan et al. (2019). Although previous theoretical works have focused on static spectral properties or eigenstate of GAA model Li et al. (2016), it remains an open question how these spectral features translate into observable non-equilibrium dynamics starting from a product state. This raises a key question for the GAA model regarding how does entanglement evolve in a system defined by the coexistence of localized and extended modes? Does it obey a volume law due to the extended states, an area law governed by the localized ones, or does a new, intermediate behavior emerge?
While EE quantifies the overall magnitude of entanglement, it does not fully resolve how quantum information propagates and where it is stored. Other tools reveal different aspects of information dynamics. For example, mutual information tracks shared correlations Wolf et al. (2008), and its extensions like tripartite mutual information can detect delocalized entanglement and information scrambling Hosur et al. (2016). Moreover, out-of-time-ordered correlators are frequently used to diagnose the spreading of local information into non-local correlations—a signature of quantum chaos Maldacena et al. (2016); Fan et al. (2017); Nahum et al. (2018). However, to obtain a more direct, spatially resolved picture of how information propagates and is retained in space-time, we employ the subsystem information capacity (SIC) Chen et al. (2025), a recently introduced probe of information propagation and retention. This quantity investigates the information flow by viewing a subsystem’s evolution as an effective quantum channel whose capacity is shaped by the system’s overall dynamics Holevo (2012). Defined as the mutual information I(A:R)I(A{:}R) between an output subsystem $A$ and a reference system $R$ initially entangled with an input region $E$, SIC measures how much of the initial quantum information can be recovered from $A$ after some time. This measure is closely related to the quantum coherent information of the effective channel $\mathcal{E}_{E\to A}$ Schumacher and Nielsen (1996); Chen et al. (2025), reflecting the system’s ability to preserve quantum information and acting as a measure of the channel’s single-shot quantum capacity Lloyd (1997); Devetak and Winter (2005). Importantly, SIC shows distinct spatial profiles in different dynamical phases—from information trapping behavior in localized systems to linear growth in extended ones Chen et al. (2025). This makes SIC an ideal diagnostic for the mixed dynamics expected in the GAA model, allowing us to ask whether the contributions of localized and extended modes can be spatially separated.
In this work, we employ two complementary quantum information probes—entanglement entropy and subsystem information capacity—to conduct a comprehensive analysis of the quench dynamics in the non-interacting GAA model. By contrasting its behavior with the abrupt localization transition in the standard AA model, we establish clear and unambiguous dynamical fingerprints of an SPME. We find that the SPME induces a crossover in both dynamical signatures, replacing the sharp transition. Specifically, the EE saturation value exhibits a persistent volume-law scaling, yet with a continuously tunable entropy controlled by the number of available extended states. Complementing this, the SIC profile directly visualizes the mixed spectral nature, interpolating between the linear ramp characteristic of extended systems and the information trapping of localized ones. Our results provide a definitive non-interacting benchmark crucial for understanding information and entanglement dynamics in interacting systems with mixed phases.
Model and Methods.— We study the dynamics of noninteracting spinless fermions on a one-dimensional lattice with open boundary conditions, described by the GAA Hamiltonian
$H=-t\sum_{i=1}^{L-1}(c_{i}^{\dagger}c_{i+1}+\text{h.c.})+\sum_{i=1}^{L}\mu_{i}\,c_{i}^{\dagger}c_{i},$ (1)
where $c_{i}^{\dagger}$ ($c_{i}$) creates (annihilates) a fermion at site $i$, $L$ is the system size, and $t$ is the nearest-neighbor hopping, set to unity ($t=1$). The on-site quasiperiodic potential is
$\mu_{i}=2\lambda\frac{\cos(2\pi bi+\phi)}{1-a\cos(2\pi bi+\phi)},$ (2)
[…]
The standard AA model ($a=0$) features a sharp localization transition at $\lambda=t$, where all single-particle eigenstates abruptly transition from being extended to localized Aubry and André (1980). For $a\neq 0$, the GAA model hosts a single-particle mobility edge at energy $E_{c}$, given by Ganeshan et al. (2015)
$aE_{c}=2\,\mathrm{sgn}(\lambda)(|t|-|\lambda|).$ (3)
This critical energy leads to the phase diagram in Fig. 1, showing a completely extended, a completely localized, and an intermediate phase with a mobility edge where extended and localized states coexist.
Figure 1: Phase diagram of the GAA model, featuring three distinct dynamical regimes based on the potential strength $\lambda/|t|$ and deformation $a$. The model has completely extended (pink) and localized (green) phases, separated by an intermediate phase (yellow) where extended and localized states coexist due to the SPME.
The key feature of the GAA model for $a\neq 0$ is the coexistence of localized and extended states. The effect of the SPME is visualized by the inverse participation ratio (IPR) of the eigenstates,
$\mathrm{IPR}_{n}=\frac{\sum_{i=1}^{L}|\psi_{n,i}|^{4}}{(\sum_{i=1}^{L}|\psi_{n,i}|^{2})^{2}},$ (4)
where $\psi_{n}$ is the $n$-th eigenstate of the system. As shown in Fig. 2 for $a=0.3$, the eigenstates are clearly separated into extended (low IPR) and localized (high IPR) modes by the energy-dependent SPME. In the following, we study the quench dynamics governed by this Hamiltonian to explore the consequences of this mixed spectral structure.
Figure 2: Energy spectrum of the GAA model for $a=0.3$ and $L=200$, with eigenstates colored by their IPR. The mobility edge $E_{c}$ (red lines), from Eq. (3), clearly separates extended states (low IPR, bright cyan) from localized states (high IPR, dark teal/black). This energy-dependent separation is a key feature of the SPME phase.
Entanglement Dynamics.— We consider unitary evolution $|\Psi(t)\rangle=e^{-iHt}|\Psi(0)\rangle$ after a quantum quench from an unentangled product state at half filling ($N=L/2$). We use the Néel state ($|1010\dots\rangle$) as the initial state throughout the main text. Results from different initial states are presented in the Supplemental Material (SM) sm (2025). To measure the dynamics, we compute the half-chain EE, $S(t)=-\mathrm{Tr}[\rho_{L/2}(t)\ln\rho_{L/2}(t)]$, where $\rho_{L/2}(t)$ is the reduced density matrix of the system’s left half.
First, we examine the early-time EE growth velocity, $v_{S}$, extracted from the initial linear growth regime (see SM sm (2025) for dynamics plots and extraction details). As shown in Fig. 3, for both the standard AA model ($a=0$) and the GAA model ($a>0$), $v_{S}$ decreases monotonically with increasing $\lambda$. Unlike the late-time saturation behavior discussed later, $v_{S}$ does not exhibit visible plateaus corresponding to the spectral gaps. This observation highlights a fundamental dynamical dichotomy in the system, wherein the early-time growth is governed by local kinetic constraints imposed by the potential strength $\lambda$, rather than the global spectral topology. Consequently, the distinct gaps in the SPME spectrum have a negligible effect on the initial spreading speed. We also note that $v_{S}$ begins to decrease immediately as $\lambda>0$, even within the fully extended phase. This confirms that any quasiperiodic potential, regardless of strength, suppresses particle propagation speed, consistent with findings in the standard AA model Longhi (2021); Shimasaki et al. (2024).
Figure 3: Early-time EE growth velocity $v_{S}$ versus $\lambda/|t|$ for different $a$. The initial state is a Néel state with $L=200$. In contrast to the saturation entanglement, $v_{S}$ shows a smooth, monotonic decrease for both the AA ($a=0$) and GAA ($a>0$) models. This indicates that early-time dynamics are primarily sensitive to the potential strength rather than the detailed gap structure of the spectrum.
Next, we analyze the long-time saturation EE, $S_{\mathrm{sat}}$, extracted by averaging the EE over a long-time window after it reaches a steady value. As shown in Fig. 4, for the standard AA model ($a=0$), $S_{\mathrm{sat}}$ drops sharply at the $\lambda=t$ transition, signaling the phase transition nature from a volume-law to an area-law phase. In contrast, the GAA models ($a>0$) exhibit a smooth crossover, indicating a gradual suppression of entanglement generation rather than a complete halt. The visible plateaus correspond to the SPME traversing the spectral gaps.
Figure 4: Saturation $E$ $S_{\mathrm{sat}}$ versus $\lambda/|t|$ after a quench from the Néel state for $L=200$. The AA model ($a=0$, blue) has a sharp drop at the $\lambda=t$ transition. The GAA models ($a>0$) show a smooth crossover, indicating partial delocalization caused by the SPME.
To rigorously probe the scaling law for the entanglement, we perform a finite-size scaling analysis of $S_{\mathrm{sat}}$ (Fig. 5) to extract the scaling exponent $\alpha$ from the relation $S_{\mathrm{sat}}\propto L^{\alpha}$, which distinguishes area-law scaling ($\alpha\approx 0$) from volume-law scaling ($\alpha\approx 1$). All cases show volume-law behavior for small $\lambda$ and area-law for large $\lambda$. Importantly, in the intermediate phase with SPME for $a\neq 0$ (e.g., $\lambda/|t|=1.0,1.3$), we find a persistent volume law, with the exponent $\alpha$ remaining approximately $1$. This confirms that the system remains partially delocalized and capable of generating extensive entanglement throughout this phase, a defining feature absent in the standard AA model where $\lambda>t$ implies full localization and an area law.
Figure 5: Scaling exponent $\alpha$ of the saturation entropy ($S_{\mathrm{sat}}\propto L^{\alpha}$) versus $\lambda/|t|$. The exponent distinguishes area-law ($\alpha\approx 0$) from volume-law ($\alpha\approx 1$) scaling. Error bars indicate the standard error of the slope obtained from the linear regression fit. The inset shows a representative linear fit of $\ln S_{\mathrm{sat}}$ versus $\ln L$ for $a=0.3$ and $\lambda/|t|=1.0$. The persistent volume law for $a>0$ at intermediate $\lambda$ confirms the delocalizing effect of the SPME.
The underlying physical mechanism is revealed in Fig. 6, which plots $S_{\mathrm{sat}}$ against the fraction of extended states, $n_{e}$ (defined as $N_{e}/L$, where $N_{e}$ counts eigenstates with $E<E_{c}$). The two quantities are strongly correlated. This striking one-to-one correspondence demonstrates that the system’s capacity to generate volume-law entanglement is directly controlled by the fraction of available extended modes, which act as resources for thermalization. In other words, the decrease in $S_{\mathrm{sat}}$ tracks the reduction in available extended modes as $\lambda$ increases.
Figure 6: Connection between saturation entanglement $S_{\mathrm{sat}}$ (blue circles, left axis) and the fraction of extended states $n_{e}$ (green triangles, right axis) versus $\lambda/|t|$ for $a=0.3$ and $L=200$. The fraction is defined as $n_{e}=N_{e}/L$, where $N_{e}$ is the number of eigenstates with energy $E<E_{c}$. The strong correlation shows that the system’s capability to develop volume-law entanglement is directly controlled by the number of available extended modes.
Figure 7: Steady-state SIC profile $I(A:R)$ versus subsystem size $|A|$ for $L=200$. The reference qubit is entangled with the center site $E=L/2$. (Top) The AA model ($a=0$) shows a sharp transition from a linear ramp to a step function. (Middle, Bottom) The GAA model ($a>0$) shows a smooth crossover. The hybrid nature of the profiles in the SPME phase—an initial jump followed by a slow ramp—is a direct visualization of the coexisting localized (information trapping) and extended (information spreading) modes.
Subsystem Information Capacity.— To move beyond the single scalar value of EE and gain a spatially resolved picture of information dynamics, we analyze the SIC Chen et al. (2025). This probe measures how quantum information, initially localized at a single site, propagates through the system. We prepare the system with a reference qubit $R$ maximally entangled with a single site $E$ of the chain
$|\Phi\rangle=\frac{1}{\sqrt{2}}(|1\rangle_{E}|0\rangle_{R}+|0\rangle_{E}|1\rangle_{R}),$ (5)
with the rest of the chain in a product state. We compute the SIC between $R$ and a subsystem $A$ of size $|A|$ centered on $E=L/2$ after the time evolution, defined as
I(A:R)=S(A)+S(R)S(AR),I(A:R)=S(A)+S(R)-S(AR), (6)
where $S(\cdot)$ is the von Neumann entropy. Note that we use base-$2$ logarithmic for the entropy convention in the SIC definition so that $I=2$ indicates the full information recovery capability. The spatial profile of steady-state $I(A:R)$ measures how information spreads, with a linear ramp signifying ballistic transport in extended systems and a step-function profile indicating information trapping in localized systems Chen et al. (2025).
The steady-state SIC profile (Fig. 7) provides a clear picture of the information distribution at late times. For the standard AA model ($a=0$), the profile shows a sharp transition from a perfect linear ramp ($\lambda/|t|<1$) to a sharp step function ($\lambda/|t|>1$). In contrast, the GAA model ($a>0$) shows a more smooth crossover. In the intermediate SPME regime, the SIC profiles are a hybrid of both behaviors, characterized by a steep rise for small ($|A|$) due to information trapping by localized modes, and a subsequent slow increase for larger ($|A|$) driven by ballistic spreading through the remaining extended modes. This interpolation between a linear ramp and information trapping is a direct and powerful dynamical signature of the SPME.
To quantify the mixed behavior, we define $\text{SIC}_{\text{jump}}$ as the SIC value for a small subsystem size (we use $|A|=5$) intended to capture the information confined by localized modes due to the trapping nature. Fig. 8 plots $\text{SIC}_{\text{jump}}$ and the fraction of localized states, $n_{l}$ (defined as $N_{l}/L$, where $N_{l}$ counts eigenstates with $E>E_{c}$). For small $\lambda$ in the fully extended phase, $\text{SIC}_{\text{jump}}$ is near zero. As $\lambda$ increases in the intermediate phase with SPME, $\text{SIC}_{\text{jump}}$ rises sharply, tracking the increase in $n_{l}$. This powerful correlation between $\text{SIC}_{\text{jump}}$ and the fraction of localized states $n_{l}$ confirms the physical origin of the mixed SIC profile, revealing that the initial jump is a direct consequence of the growing fraction of localized states, which trap a corresponding amount of information around the initial site. We also present the late-time SIC profiles when the reference site is coupled to the boundary site in the SM sm (2025), which further corroborates our findings.
Figure 8: Connection between the initial SIC jump, $\text{SIC}_{\text{jump}}$ (blue circles, left axis), and the fraction of localized states $n_{l}$ (green triangles, right axis) for the GAA model with $a=0.3$ and $L=200$. The fraction is defined as $n_{l}=N_{l}/L$, where $N_{l}$ is the number of eigenstates with energy $E>E_{c}$. The close connection shows that the emergence of information confinement, captured in $\text{SIC}_{\text{jump}}$, is directly caused by the increase in localized states.
Table 1: Comparison of dynamical signatures in the standard AA and GAA models within SPME phases.
Probe Standard AA GAA
($a=0$) ($a>0$)
$E$ ($S_{\text{sat}}$) Sharp transition: SPME phase:
Volume $\to$ Area Persistent volume-law
SIC Profile Sharp transition: SPME phase:
ramp $\to$ step-func. Hybrid jump+ramp
Discussions and conclusion.—Our comprehensive analysis reveals that the SPME fundamentally reshapes the system’s dynamical response. Our work goes beyond the qualitative expectation of a “smoothed” transition to uncover the precise dynamical mechanisms governing this crossover. Our results show clear differences between the standard AA model ($a=0$) and the GAA model ($a\neq 0$) with new physical insights. In the AA model ($a=0$), the transition is sudden, with $S_{\mathrm{sat}}$ switches from volume-law to area-law scaling abruptly, and the SIC profile changes from a ramp to a step. In contrast, the GAA model ($a\neq 0$) shows a smooth crossover. This is characterized by intermediate $S_{\mathrm{sat}}$ values, mixed SIC profiles, and a persistent volume law with a decreasing entropy. This smoothing is a direct result of the SPME, which allows localized and extended modes to coexist and control the balance between information spreading and confinement. Our main findings are summarized in Table 1. Our results provide clear non-interacting dynamical signatures of a mobility edge, establishing an important reference for understanding thermalization in more complex quasiperiodic systems. These insights are relevant to the current debate on MBL in the presence of an SPME Li et al. (2015); Modak and Mukerjee (2015); Deng et al. (2017); Huang et al. (2023). While we focus on the paradigmatic GAA model, our findings on the dynamical signatures of coexisting states are expected to provide insights for other systems hosting mobility edges, such as those with long-range interactions or in higher dimensions Li et al. (2016); Zhang and Yao (2018); Huang et al. (2023).
In conclusion, we have studied the quantum dynamics in the generalized Aubry-André model, a paradigmatic system with a single-particle mobility edge. By leveraging the complementary power of entanglement entropy and subsystem information capacity, we have demonstrated that the SPME fundamentally reshapes the system’s dynamical response, converting the sharp localization transition into a smooth crossover. Instead of the sudden transition in the standard AA model, the SPME facilitates a regime with smoothly varying entanglement, hybrid information-spreading profiles, and a robust volume law with tunable entropy. Our findings highlight that these quantum information probes are highly sensitive to a system’s spectral structure and that mobility edges reshape dynamical phase boundaries. This work provides a clear and crucial non-interacting reference for localization dynamics in quasiperiodic systems and paves the way for future studies of interacting systems, where the interplay between SPMEs and many-body effects remains a major open question. Furthermore, understanding the control over information flow offered by an SPME, from ballistic transport to perfect trapping, holds potential relevance for quantum technologies.
Acknowledgements. S.X.Z is supported by Quantum Science and Technology-National Science and Technology Major Project (No. 2024ZD0301700) and the National Natural Science Foundation of China (No. 12574546). Y.Q.C is supported by National Natural Science Foundation of China (No. 12504599) and NSAF (No. U2330401).
Data availability. Code implementation and numerical data for this manuscript are publicly accessible in Ref.  [^1] https://gitee.com/yuqiqing/sic_in_-gaa_model.